OpenStudy (anonymous):

A ball was thrown at 50miles per second and it was to reach 3000miles what is the angle that it needs to be thrown to hit the destination? with the formula r=(1/32)(Vo)^2sin2

OpenStudy (lgbasallote):

where r is...what?

OpenStudy (anonymous):

What?

OpenStudy (lgbasallote):

you gave the formula r=(1/32)(Vo)^2sin2 what does r represent?

OpenStudy (anonymous):

The angle I think

OpenStudy (anonymous):

No, it's 3000 miles...I'm pretty sure

OpenStudy (lgbasallote):

so where does the angle come in?

OpenStudy (anonymous):

This is what I got Vo=50m/s 50*5280=26400ft/s r=3000 miles=15840000 ft r=(1/32)(Vo)^2sin2t 15840000*32= (26400)2 sin2t 0.7273=sin2t 2t=46.6584 t= 23.3292 degrees

OpenStudy (lgbasallote):

is it \[\large r = \frac{1}{32} (V_o)^2\sin 2\] or \[\Large r = \frac{1}{32} (v_o)^2 \sin \theta\]

OpenStudy (anonymous):

THe second one

OpenStudy (lgbasallote):

why did you multiply 5280 to 50 again?

OpenStudy (swissgirl):

so u gotta work in ft????

OpenStudy (swissgirl):

everything is in miles neways

OpenStudy (anonymous):

That's what I'm thinking...

OpenStudy (swissgirl):

ok well either way u shld get the same answer lets try it out in miles

OpenStudy (anonymous):

Alright...

OpenStudy (swissgirl):

nooo u r right for some reason u gotta work in ft

OpenStudy (anonymous):

not this again...

OpenStudy (swissgirl):

hahahha

OpenStudy (anonymous):

Am I right?

OpenStudy (swissgirl):

Like if u are working in ft that is correct. U got that answer from yahoo answers :P

OpenStudy (swissgirl):

Ya i just did it like a bunch of times and like I am figuring u have to work in ft so ya there is ur answer

OpenStudy (anonymous):

What's yahoo answers?